a 12-kg block is acted upon by two forces that are in opposite direction, one a 540 Newton and the other of 1260 N. Determine the acceleration produced. How is the direction of the acceleration related to the forces


1
Expert's answer
2021-02-28T07:30:12-0500

According to Newton's second law,

"F_{net}=ma"

"F_1-F_2=ma"

The acceleration of the block is

"a=\frac{F_1-F_2}{m}"

"=\frac{1260 N-540 N}{12 kg}"

="60 m\/s^2"

Direction is towards large force.